Mathch one of the following bonds with each of the categories shown below

Match a bond order with each category given. Matcand explain thank you
Answer #1

More the bond order more will be the bond strength, lesser will be the bond length and thus more energy will be required to break the bond

Thus, smallest bond order = A-B

Greatest bond order = A(triple bond)B

Smallest bond energy = A-B

Greatest bond energy = A(triple bond)B

Shortest bond length = A(triple bond)B

Greatest bond length = A-B